OpenStudy (anonymous):

Find the equation of the line with the given properties. express the equation in general form or slope-intercept form. Perpendicular to the line 5x+y=-14; contains the point (-5,10)

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

y=-5x-14?

OpenStudy (ipwnbunnies):

Yes, that's the equation of the given line.

OpenStudy (ipwnbunnies):

The slope of the perpendicular line is: \[m(perpendicular) = \frac{ -1 }{ m }\]

OpenStudy (ipwnbunnies):

Then, use point-slope form to get the equation of the line perpendicular to the given line.

OpenStudy (anonymous):

see i think thats why i was confused about point slope form lol

OpenStudy (anonymous):

y=-1/5x+9 ?

OpenStudy (anonymous):

wait...no.. the two negatives in the fraction cancel out right?

OpenStudy (ipwnbunnies):

Right, the slope of the perp. line should be (1/5)

OpenStudy (anonymous):

so it would be y=1/5x+11

OpenStudy (ipwnbunnies):

(y-10) = (1/5)(x + 5) Yes, you're right. ^ I was writing it out for myself lol.

OpenStudy (anonymous):

YAY! thanks! =)